# Convolution goes higher-order: a biologically inspired mechanism empowers image classification

### v1

@@ Abstract
We propose a novel approach to image classification inspired by complex nonlinear biological visual processing, whereby classical convolutional neural networks (CNNs) are equipped with learnable higher-order convolutions. Our model incorporates a Volterra-like expansion of the convolution operator, capturing multiplicative interactions akin to those observed in early and advanced stages of biological visual processing. We evaluated this approach on synthetic datasets by measuring sensitivity to testing higher-order correlations and performance in standard benchmarks (MNIST, FashionMNIST, CIFAR10, CIFAR100 and Imagenette). Our architecture outperforms traditional CNN baselines, and achieves optimal performance with expansions up to 3rd/4th order, aligning remarkably well with the distribution of pixel intensities in natural images. Through systematic perturbation analysis, we validate this alignment by isolating the contributions of specific image statistics to model performance, demonstrating how different orders of convolution process distinct aspects of visual information. Furthermore, Representational Similarity Analysis reveals distinct geometries across network layers, indicating qualitatively different modes of visual information processing. Our work bridges neuroscience and deep learning, offering a path towards more effective, biologically inspired computer vision models. It provides insights into visual information processing and lays the groundwork for neural networks that better capture complex visual patterns, particularly in resource-constrained scenarios.
@@ 1 Introduction
”The correlation over space translates into a correlation over order.” (Koenderink & van Doorn, 2018)
@@
Convolutional Neural Networks (CNNs) have become the cornerstone of modern computer vision tasks, demonstrating remarkable performance across a wide range of applications (Krizhevsky et al., 2012), (Simonyan & Zisserman, 2014), (He et al., 2016). The success of CNNs is largely attributed to their ability to capture local patterns and hierarchical features in images through their layered structure and weight sharing mechanism (LeCun et al., 2015), (Zeiler & Fergus, 2014).
@@
However, natural images contain complex correlations and higher-order statistics that extend beyond simple linear relationships (Field, 1993), (Olshausen & Field, 1996), (Schwartz & Simoncelli, 2001). These include texture patterns, edge interactions, and intricate geometric structures that are prevalent in real-world visual scenes. Importantly, these correlations exhibit a hierarchical structure, with higher-order correlations becoming increasingly sparse (Koenderink & van Doorn, 2018). Traditional CNNs, particularly those with limited depth, often struggle to effectively exploit these higher-order correlations (Gatys et al., 2016), (Ustyuzhaninov et al., 2022), (Ding et al., 2020). The limitation of standard CNNs in capturing higher-order correlations stems from their reliance on linear convolutions followed by pointwise nonlinearities (Yu & Salzmann, 2017). Although deeper networks can approximate complex functions through the composition of many layers, this approach may be computationally expensive and require large amounts of training data.
To address this limitation, we propose a learnable extension of the classical linear convolution operator to incorporate higher-order terms, akin to a Volterra expansion (Volterra, 1887), (Volterra, 1930). This approach allows for the direct modeling of multiplicative interactions between neighboring pixels, enabling the network to capture complex local structures more effectively, even in shallow architectures.
@@
We design models by incorporating higher-order convolutions, and compare their performance against standard CNN baselines. We evaluated our model on a range of datasets, beginning with synthetic images typically used to assess sensitivity to higher-order correlations. Subsequently, we extended our experiments to compare our model against CNN baselines on widely-used datasets including MNIST, FashionMNIST, CIFAR10, CIFAR100 and Imagenette. Our findings reveal that these higher-order convolutions improve image classification performance, offering a promising direction for advancing deep learning models in computer vision tasks.
@@
Interestingly, our results show optimal performance at the 3rd or 4th order, aligning remarkably with the distribution of pixel intensities in natural images as analyzed by Koenderink & van Doorn (2018) and Tkačik et al. (2010). This alignment suggests that our approach effectively captures the fundamental structure of visual information in the natural world. Through systematic perturbation analysis, we isolate the contributions of specific image statistics to model performance, providing insight into how different orders of convolution process distinct aspects of visual information. Furthermore, we employ Representational Similarity Analysis (RSA) (Kriegeskorte et al., 2008) to investigate the internal representations learned by our model compared with standard CNNs, revealing distinct geometries that indicate different strategies for encoding and processing visual information.
@@
Our approach addresses the limitations of simple pointwise nonlinearities in capturing local image structure, especially in networks of limited depth or width. In particular, we show that incorporating such computations at the earliest stages of image processing, even within the first layer of a neural network, yields significant benefits. This finding suggests that models benefit from embedding these computations via direct engagement with the superficial structure of images (Koenderink, 1984).
@@ 1.1 Biological Inspiration and Related Work
Our work draws inspiration from the sophisticated mechanisms of biological visual systems, where visual information processing extends beyond simple pointwise nonlinearities and first-order convolutions (Fitzgerald & Clark, 2015; Clark et al., 2014; Lettvin et al., 1959; Peterhans & Von der Heydt, 1993; Levitt et al., 1994). These non-pointwise nonlinearities, which consider spatial relationships between multiple input points, enable more effective extraction and integration of spatial information (Krieger et al., 1997; Zetzsche & Barth, 1990; Barth et al., 1993; Koenderink & Richards, 1988).
This complex processing begins in the retina, where certain cells exhibit responses to visual stimuli that cannot be explained by simple pointwise nonlinearities (Lettvin et al., 1959). In the visual cortex, specialized neurons respond to complex visual features like corners or line endings (Hubel & Wiesel, 1965), performing advanced computations that integrate information across their receptive fields. Such multidimensional, non-pointwise operators are fundamental to visual processing, as emphasized by Zetzsche et al. (1993) and Koenderink & Richards (1988). Similar mechanisms appear across species - fruit flies employ non-pointwise computations for motion detection as early as the second dendritic layer (Fitzgerald & Clark, 2015), suggesting these computations are fundamental for efficient visual processing (Gilbert, 2007). Recent research has further demonstrated that single dendrites can solve complex problems like XOR (Xu et al., 2012; Ran et al., 2020), a capability that pointwise nonlinearities struggle to model (Minsky & Papert, 1969).
In computational research, the concept of nonlinear receptive fields has been explored through various approaches. Berkes & Wiskott (2006) analyzed inhomogeneous quadratic forms as receptive fields, while Zoumpourlis et al. (2017) extended this to CNN-based learning. For video classification, Roheda et al. (2024) introduced Volterra Neural Networks using a low-rank approach. Related developments in fine-grained image classification include bilinear CNN models (Lin et al., 2015) and compact bilinear pooling (Gao et al., 2016), with applications extending to computational neuroscience (Ahrens et al., 2008).
Our approach differs fundamentally from previous methods by embedding higher-order operations within the convolutional operator itself, rather than applying them after feature extraction. This design choice enables the network to learn and apply complex nonlinear transformations throughout its entire depth, creating a more flexible architecture that better constitutes an algorithmic implementation of biological visual processing. By integrating these operations directly into the network’s core building blocks, we enable more efficient feature learning while maintaining architectural simplicity.
@@ 1.2 Beyond Pointwise Nonlinearities
$=\alpha_{0}+\color[rgb]{1,0,0}{\alpha_{1}}\color[rgb]{0,0,1}{w_{1}}\color[rgb]{0,0,0}{x_{1}}+\color[rgb]{1,0,0}{\alpha_{1}}\color[rgb]{0,1,0}{w_{2}}\color[rgb]{0,0,0}{x_{2}}+\color[rgb]{1,.5,0}{\alpha_{2}}\color[rgb]{0,0,1}{w_{1}}\color[rgb]{0,0,0}{{}^{2}x_{1}}^{2}+\color[rgb]{1,.5,0}{\alpha_{2}}\color[rgb]{0,1,0}{w_{2}}\color[rgb]{0,0,0}{{}^{2}x_{2}}^{2}+2\color[rgb]{1,.5,0}{\alpha_{2}}\color[rgb]{0,0,1}{w_{1}}\color[rgb]{0,1,0}{w_{2}}\color[rgb]{0,0,0}{x_{1}}\color[rgb]{0,0,0}{x_{2}}+\ldots$
This tied-weight issue becomes particularly problematic when the network is not deep enough or wide enough to compensate, as the precise bounds for enough depth or width depend on the universal approximation theorem and are often challenging to determine exactly (Bahri et al., 2020). Consequently, the standard pointwise nonlinearity with linear summation may fail to capture complex relationships in the data, especially in shallower or narrower networks.
To address this limitation, we propose a novel approach that can be viewed as an implementation of non-pointwise nonlinearities. Our method, which will be described in detail in Section 2, can be seen as a learnable generalized 2-D detector model (Zetzsche & Barth, 1990) representing a receptive field that responds to multiplicative interactions between inputs, similar to how some biological neurons process visual information (see Subsection 1.1). This approach offers a solution for capturing complex correlations at the level of a single mechanism, allowing for more flexible and powerful representations even in compact network architectures.
2 Model Architecture: Higher-Order Convolution
To perform more complex computations while preserving the benefits of locality and weight sharing that are inherent in standard convolutions, we propose the concept of higher-order convolution. This approach extends the traditional convolution operator to include higher-order terms, allowing for more sophisticated feature extraction, and resulting in a stackable layer for empowering CNNs.
In order to explain our higher-order convolutional layer, we start by considering an input image patch P with n elements, reshaped as a vector x. Standard convolution can be expressed in terms of linear filtering, relating input x and output y as follows: $y(\textbf{x})=b+\sum_{i=1}^{n}w_{1}^{i}x_{i}$ where $w^{i}_{1},i=1,...,n$ represents the weights of the convolution kernel.
We expand this function to include quadratic, cubic, and higher-order terms. The general form of this expansion is:
$y(\textbf{x})=b+\sum_{i=1}^{n}w_{1}^{i}x_{i}+\sum_{j=1}^{n}\sum_{i=1}^{n}w_{2}^{ij}x_{i}x_{j}+\sum_{k=1}^{n}\sum_{j=1}^{n}\sum_{i=1}^{n}w_{3}^{ijk}x_{i}x_{j}x_{k}+...$ (2)
@@ 2 Model Architecture: Higher-Order Convolution
$Y_{l,m}=b$ $+\sum_{i,j}w_{1}^{ij}X_{l+i,m+j}$
$+\sum_{i,j}\sum_{k,h}w_{2}^{ijkh}X_{l+i,m+j}X_{l+k,m+h}$
$+\sum_{i,j}\sum_{k,h}\sum_{p,q}w_{3}^{ijkhpq}X_{l+i,m+j}X_{l+k,m+h}X_{l+p,m+q}$
+$\ldots$+ … (3)
@@
By symmetry considerations, the total number of parameters in this expansion grows as: $n_{V}=\frac{(n+p)!}{n!p!}$ where $p$ represents the order of the term and $n$ the size of the kernel. To address potential issues with the relative magnitudes of higher-order terms, we implement a normalization strategy. For kernels of order greater than the 1st, we apply a scaling factor $s$ calculated as $s=\frac{1}{\sqrt{n_{V}}}$. The scaling factor helps to balance the contribution of higher-order terms relative to the first-order (linear) terms, promoting stability and preventing higher-order terms from dominating network behavior during training.
Figure 1 A illustrates this concept for binary image, showing how the classical convolution is extended to include higher-order terms for a single patch. In practice, this operation is applied across the entire image (as shown in the equation above), resulting in a feature map for each order of the expansion, which are then summed together before a standard pointwise nonlinearity as ReLU (see for example, Figure 3 B).
Figure 1: Extending classical convolution (A) A toy example of how higher-order convolution is implemented. (Left) We consider a patch located at the bottom-left of the input image, and by flattening it to a vector. (Right) Classical convolution would correspond to the 1st order term, where a set of weights (the kernel) are learnt and then translated onto other patches (weight-sharing). We extend classical convolution by expanding the first term: for the 2nd order, we compute the outer product of the original vector x with itself (thus obtaining a matrix) and the corresponding set of weights; similarly for the 3rd order, we iterate the same operation one more time and learn the associated set of weights. Feature maps of the three convolutions are then summed up before a standard pointwise nonlinearity as, for example, a ReLU function. (B) Cumulative explained variance of principal components for standard CNN model vs Higher-order CNN models with 2nd and 3rd expansion quantitatively confirms the tied-weight issue for classical models.
At this point, we have outlined the construction of a higher-order convolutional layer. This layer can functionally replace a standard convolutional layer in neural network architectures and is compatible with training through backpropagation.
@@ 2.1 A Quantitative Perspective on the Tied-Weight Issue
To demonstrate how a Higher-order CNN (HoCNN) addresses the tied-weight issue introduced in Section 1.2, we conducted a systematic analysis of weight independence in the network’s representations. We generated a fixed input image (32 × 32 pixels) containing binary textures exhibiting all possible 1-, 2-, 3-, and 4-point correlations for 2×2 patches, corresponding to fixing the $x_{i}$ terms in Equation 1. We then randomly initialized the same model architecture 10,000 times and analyzed the activations after the convolutional block (Conv/HoConv + BatchNorm + ReLU + Max Pooling) for two architectures: a CNN with 10 kernels of size 2×2 and a HoCNN with 2 kernels of size 2×2 with higher-order expansions.
To quantify the degree of weight independence, we employed Principal Component Analysis of these activations, measuring the number of principal components (PCs) needed to explain 95% of the variance. Our hypothesis was that tied weights would reduce the number of principal components, while greater weight independence would require more components to capture the same variance. The results confirmed our expectations (see Figure 1 B): the CNN required only 0.9% of PCs (87 out of 9610) to explain 95% of the variance, while the HoCNN required 5.3% of PCs (102 out of 1922) with second-order expansion and 8.3% (159 out of 1922) with third-order expansion. While the CNN has more total components due to its higher channel count (5×), additional analyses with matched channel counts yielded similar results.
These findings quantitatively demonstrate that higher-order convolutions introduce more independent weights, leading to a richer representation space with reduced weight coupling. The increasing number of components needed for variance explanation directly shows how our approach mitigates the tied-weight issue inherent in standard CNNs. Additional analyses with various nonlinearities showed consistent results, further supporting these conclusions (see Appendix A.5).

### v2

@@ Abstract
We propose a novel enhancement to Convolutional Neural Networks (CNNs) by incorporating learnable higher-order convolutions inspired by nonlinear biological visual processing. Our model extends the classical convolution operator using a Volterra-like expansion to capture multiplicative interactions observed in biological vision. Through extensive evaluation on standard benchmarks and synthetic datasets, we demonstrate that our architecture consistently outperforms traditional CNN baselines, achieving optimal performance with 3rd/4th order expansions. Systematic perturbation analysis and Representational Similarity Analysis reveal that different orders of convolution process distinct aspects of visual information, aligning with the statistical properties of natural images. This biologically-inspired approach offers both improved performance and deeper insights into visual information processing.
@@ 1 Introduction
"The correlation over space translates into a correlation over order." Koenderink and van Doorn [1]
@@
Convolutional Neural Networks (CNNs) have become the cornerstone of modern computer vision tasks, demonstrating remarkable performance across a wide range of applications [2], [3], [4]. The success of CNNs is largely attributed to their ability to capture local patterns and hierarchical features in images through their layered structure and weight sharing mechanism [5], [6].
@@
However, natural images contain complex correlations and higher-order statistics that extend beyond simple linear relationships [7], [8], [9]. These include texture patterns, edge interactions, and intricate geometric structures that are prevalent in real-world visual scenes. Importantly, these correlations exhibit a hierarchical structure, with higher-order correlations becoming increasingly sparse [1]. Traditional CNNs, particularly those with limited depth, often struggle to effectively exploit these higher-order correlations [10], [11], [12]. The limitation of standard CNNs in capturing higher-order correlations stems from their reliance on linear convolutions followed by pointwise nonlinearities [13]. Although deeper networks can approximate complex functions through the composition of many layers, this approach may be computationally expensive and require large amounts of training data. To address this limitation, we propose a learnable extension of the classical linear convolution operator to incorporate higher-order terms, akin to a Volterra expansion [14], [15]. This approach allows for the direct modeling of multiplicative interactions between neighboring pixels, enabling the network to capture complex local structures more effectively, even in shallow architectures.
@@
We design models by incorporating higher-order convolutions, and compare their performance against standard CNN baselines. We evaluated our models on a range of datasets, beginning with synthetic images typically used to assess sensitivity to higher-order correlations. Subsequently, we extended our experiments to compare our model against CNN baselines on widely-used datasets including MNIST, FashionMNIST, CIFAR10, CIFAR100, Imagenette and Imagenet. Our findings reveal that these higher-order convolutions improve image classification performance, offering a promising direction for advancing deep learning models in computer vision tasks.
@@
Interestingly, our results show optimal performance at the 3rd or 4th order, aligning remarkably with the distribution of pixel intensities in natural images as analyzed by Koenderink and van Doorn [1] and Tkačik et al. [16]. This alignment suggests that our approach effectively captures the fundamental structure of visual information in the natural world. Through systematic perturbation analysis, we isolate the contributions of specific image statistics to model performance, providing insight into how different orders of convolution process distinct aspects of visual information. Furthermore, we employ Representational Similarity Analysis (RSA) [17] to investigate the internal representations learned by our model compared with standard CNNs, revealing distinct geometries that indicate different strategies for encoding and processing visual information.
@@
Our approach addresses the limitations of simple pointwise nonlinearities in capturing local image structure, especially in networks of limited depth or width. In particular, we show that incorporating such computations at the earliest stages of image processing, even within the first layer of a neural network, yields significant benefits. This finding suggests that models benefit from embedding these computations via direct engagement with the superficial structure of images [18].
@@ 1.1 Biological Inspiration and Related Work
Our work draws inspiration from biological visual systems, where processing extends beyond simple pointwise nonlinearities and first-order convolutions [19, 20, 21, 22, 23]. These non-pointwise nonlinearities enable more effective extraction and integration of spatial information [24, 25, 26, 27], beginning in the retina [21] and continuing in the visual cortex where specialized neurons respond to complex features like corners or line endings [28].
In computational research, the concept of nonlinear receptive fields has been explored through various approaches. Berkes and Wiskott [29] analyzed inhomogeneous quadratic forms as receptive fields, while Li et al. [30] provided theoretical insights on Volterra-based convolutions in neural networks. Zoumpourlis et al. [31] implemented Volterra-inspired convolutions primarily in the first layer of a Wide ResNet-28, but failed to demonstrate significant performance improvements and limited their analysis to second-order expansions. In contrast, our work extends this foundation by implementing and systematically analyzing expansions up to the 4th order, where we demonstrate that 3rd/4th order expansions yield optimal performance—a finding we uniquely connect to the statistical properties of natural images documented by Koenderink and van Doorn [1].
For video classification, Roheda et al. [32] introduced Volterra Neural Networks using cascaded 2nd-order filters with low-rank approximations for spatiotemporal tasks, creating separate spatial and temporal streams that are later fused. Our approach differs fundamentally, as we focus on static image classification with direct implementation of higher-order terms (up to 4th order) within the convolution operation itself, preserving parameter independence across orders for more flexible feature learning. While Roheda et al. [32] emphasize computational efficiency through approximations, our work provides deeper insights through systematic perturbation analysis and representational similarity analysis that reveals how different orders process distinct aspects of visual information. Related developments in fine-grained image classification include bilinear CNN models [33] and compact bilinear pooling [34], with applications extending to computational neuroscience [35].
Our approach extends and complements these previous higher-order methods by embedding higher-order operations directly within the convolutional operator, creating independent feature maps for each order that are combined before nonlinearities. This design choice enables the network to learn and apply complex nonlinear transformations that better constitute an algorithmic implementation of biological visual processing. Through our structured perturbation analysis, representation geometry analysis, and controlled texture experiments, we provide novel insights that go beyond performance improvements, demonstrating why and how higher-order processing benefits visual representation.
@@ 1.2 Beyond Pointwise Nonlinearities
$=\alpha_{0}+\color[rgb]{1,0,0}{\alpha_{1}}\color[rgb]{0,0,1}{w_{1}}\color[rgb]{0,0,0}{x_{1}}+\color[rgb]{1,0,0}{\alpha_{1}}\color[rgb]{0,1,0}{w_{2}}\color[rgb]{0,0,0}{x_{2}}+\quad\color[rgb]{1,.5,0}{\alpha_{2}}\color[rgb]{0,0,1}{w_{1}}\color[rgb]{0,0,0}{{}^{2}x_{1}^{2}}+\color[rgb]{1,.5,0}{\alpha_{2}}\color[rgb]{0,1,0}{w_{2}}\color[rgb]{0,0,0}{{}^{2}x_{2}^{2}}+2\color[rgb]{1,.5,0}{\alpha_{2}}\color[rgb]{0,0,1}{w_{1}}\color[rgb]{0,1,0}{w_{2}}\color[rgb]{0,0,0}{x_{1}x_{2}}+\ldots$
This tied-weight issue becomes particularly problematic when the network is not deep enough or wide enough to compensate, as the precise bounds for enough depth or width depend on the universal approximation theorem and are often challenging to determine exactly [36]. Consequently, the standard pointwise nonlinearity with linear summation may fail to capture complex relationships in the data, especially in shallower or narrower networks. To address this limitation, we propose a novel approach that can be viewed as an implementation of non-pointwise nonlinearities. Our method, which will be described in detail in Section 2, can be seen as a learnable generalized 2-D detector model [25] representing a receptive field that responds to multiplicative interactions between inputs, similar to how some biological neurons process visual information (see Subsection 1.1). This approach offers a solution for capturing complex correlations at the level of a single mechanism, allowing for more flexible and powerful representations even in compact network architectures.
2 Higher-Order Convolution
To perform more complex computations while preserving the benefits of locality and weight sharing that are inherent in standard convolutions, we propose the concept of higher-order convolution. This approach extends the traditional convolution operator to include higher-order terms, allowing for more sophisticated feature extraction.
In order to explain our higher-order convolutional layer, we start by considering an input image patch P with n elements, reshaped as a vector x. Standard convolution can be expressed in terms of linear filtering, relating input x and output y as follows: $y(\textbf{x})=b+\sum_{i=1}^{n}w_{1}^{i}x_{i}$ where $w^{i}_{1},i=1,...,n$ represents the weights of the convolution kernel. We expand this function to include quadratic, cubic, and higher-order terms. The general form of this expansion is:
$y(\textbf{x})=b+\sum_{i=1}^{n}w_{1}^{i}x_{i}+\sum_{j=1}^{n}\sum_{i=1}^{n}w_{2}^{ij}x_{i}x_{j}+\sum_{k=1}^{n}\sum_{j=1}^{n}\sum_{i=1}^{n}w_{3}^{ijk}x_{i}x_{j}x_{k}+$...
@@ 2 Higher-Order Convolution
$Y_{l,m}=b+\sum_{i,j}w_{1}^{ij}X_{l+i,m+j}+\sum_{i,j}\sum_{k,h}w_{2}^{ijkh}X_{l+i,m+j}X_{l+k,m+h}$
$+\sum_{i,j}\sum_{k,h}\sum_{p,q}w_{3}^{ijkhpq}X_{l+i,m+j}X_{l+k,m+h}X_{l+p,m+q}+\ldots$ (2)
@@
By symmetry considerations, the total number of parameters in this expansion grows as: $n_{V}=\frac{(n+p)!}{n!p!}$ where $p$ represents the order of the term and $n$ the size of the kernel. To address potential issues with the relative magnitudes of higher-order terms, we implement a normalization strategy. For kernels of order greater than the 1st, we apply a scaling factor $s$ calculated as $s=\frac{1}{\sqrt{n_{V}}}$. The scaling factor helps to balance the contribution of higher-order terms relative to the first-order (linear) terms, promoting stability and preventing higher-order terms from dominating network behavior during training. Figure 1 A illustrates this concept for binary image, showing how the classical convolution is extended to include higher-order terms for a single patch. In practice, this operation is applied across the entire image, resulting in a feature map for each order of the expansion, which are then summed together before a standard pointwise nonlinearity as ReLU (see for example, Figure 2 B). This layer can functionally replace a standard convolutional layer in neural network architectures and is compatible with training through backpropagation.
Figure 1: Extending classical convolution (A) Implementation of higher-order convolution: Input patch is flattened to vector x. 1st order uses classical convolution with weight-sharing. 2nd and 3rd orders compute outer products - respectively one and two times - of x with itself and dot product with the corresponding weights. Feature maps are then summed before ReLU activation. (B) Cumulative explained variance of principal components for standard CNN model vs Higher-order CNN models with 2nd and 3rd expansion quantitatively confirms the tied-weight issue for classical models. Inset: Example of the 32×32 binary texture image used in this analysis, containing all possible 1-, 2-, 3-, and 4-point correlations for 2×2 patches.
@@ 2.1 A Quantitative Perspective on the Tied-Weight Issue
To demonstrate how a Higher-order CNN (HoCNN) addresses the tied-weight issue introduced in Section 1.2, we conducted a systematic analysis of weight independence in the network’s representations. We generated a fixed input image (32 × 32 pixels) containing binary textures exhibiting all possible 1-, 2-, 3-, and 4-point correlations for 2×2 patches (see inset in Figure 1B), corresponding to fixing the $x_{i}$ terms in Equation 1. We then randomly initialized the same model architecture 10,000 times and analyzed the activations after the convolutional block (Conv/HoConv + BatchNorm + ReLU + Max Pooling) for two architectures: a CNN with 10 kernels of size 2×2 and a HoCNN with 2 kernels of size 2×2 with higher-order expansions.
To quantify the degree of weight independence, we employed Principal Component Analysis of these activations, measuring the number of principal components (PCs) needed to explain 95% of the variance. Our hypothesis was that tied weights would reduce the number of principal components, while greater weight independence would require more components to capture the same variance. The results confirmed our expectations (see Figure 1 B): the CNN required only 0.9% of PCs (87 out of 9610) to explain 95% of the variance, while the HoCNN required 5.3% of PCs (102 out of 1922) with second-order expansion and 8.3% (159 out of 1922) with third-order expansion. While the CNN has more total components due to its higher channel count (5×), additional analyses with matched channel counts yielded similar results. These findings quantitatively demonstrate that higher-order convolutions introduce more independent weights, leading to a richer representation space with reduced weight coupling. The increasing number of components needed for variance explanation directly shows how our approach mitigates the tied-weight issue inherent in standard CNNs. Additional analyses with various nonlinearities showed consistent results, further supporting these conclusions (see Appendix A.11).